how do you solve equations by adding or subtracting fractions with variables? Example:m-(-7/10)= -1 1/5
Answer by Alan3354(69443)

m-(-7/10)= -1 1/5
m + (7/10) = -6/5
Eliminate the fractions
50m + 35 = -60
50m = -95
m = -19/10